PHOTODISSOCIATION OF NO ISOLATED IN SOLID PARAHYDROGEN

\underline{DAVID T. ANDERSON}, LEIF O. PAULSON and SHARON C. KETTWICH, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

The in situ photochemistry of dopant molecules isolated in solid parahydrogen (pH_{2}) typically differs from analogous studies in rare gas crystals for two main reasons: (1) solid pH_{2} has a negligible cage effect so that photodissociation of a precursor molecule can lead efficiently to well-separated fragments, and (2) radical fragments can potentially react with the pH_{2} matrix. Our group is currently studying the 193 nm photochemistry of a number of precursor molecules isolated in solid pH_{2} via high-resolution FTIR spectroscopy in hopes of identifying trapped atomic species. Currently we are revisiting the photochemistry of NO in solid pH_{2} that has previously been investigated by Momose and co-workers.^{a} The 193 nm photodynamics of NO in rare gas matrices has also been extensively studied.^{b} The studies by Momose showed that NO is photolabile at 193 nm and that the product N and O atoms react to form NH_{3} and H_{2}O, respectively. Preliminary experiments in this laboratory show evidence for the production of the imidogen (NH) radical as well, and the most recent results and analysis will be presented.

aM. Fushitani and T. Momose, Low Temp. Phys. 29, 740 (2003).

bJ. Eloranta, K. Vaskonen, H. Hakkanen, T. Kiljunen, and H. Kunttu, J. Chem. Phys. 109, 7784 (1998).